OpenStudy (anonymous):

if you have (a*x^4)*(b*x^6), does that =ab(x)^10? so, in other words, if I had... 10(x^7)*[(10x^3)+(5x^2)-7] how would you simplify it?

OpenStudy (unklerhaukus):

yes \[(a \times x^4) \times (b\times x^6)=ab(x)^{10}\] and \[10x^7 \times [10x^3+5x^2-7]=100x^{21}+50x^9+70x^7 \]